1. Simplify each expression, for n = 5 a) 8n + 2 b) n2 - 2n+n d) 19 - 3n + e) (3n - 3) + 15-n-18
Answer #1

@ 8nts n=5 = 8x5+5 - Yots = 45 1 © 222h thnos -32-h = (5) ²5 = 25-5-20 19-3n+25 , n=5 > 19-3x5+255 - 19-15+5 = 19-10 (30-3) +